# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v3

@@ VI Simultaneous symmetry breaking in spontaneous Floquet states
of the constrained Lagrangian (20), where we recall that $\{K,\mathbf{Q}\}$ are still a set of conserved charges. The resulting equation of motion is
@@
When inserted in the usual time-dependent GP equation (22), taking $\Psi(\mathbf{x},0)=\Psi_{0}(\mathbf{x},0)$ yields
@@
$[M_{0}(t)-i\partial_{t}]z_{\alpha}=0,~{}[M_{0}(t)-i\partial_{t}]z_{A}=i(\partial_{A}\lambda^{\beta})z_{\beta},$ (120)
@@
$[M_{0}(t)-i\partial_{t}]z_{\varepsilon,\nu}=\varepsilon z_{\varepsilon,\nu},$ (121)
with $z_{\varepsilon,\nu}(\mathbf{x},t+T)=z_{\varepsilon,\nu}(\mathbf{x},t)$, $\varepsilon$ the quasi-energy (defined modulo $\omega=2\pi/T$), and $\nu$ a discrete index labeling the solution. With the help from the conventional Floquet theory and from the results of Sec. III.2, it is shown that the orthogonality relations (57) still hold:
$(z_{\varepsilon,\nu}|z_{\varepsilon^{\prime},\nu^{\prime}})=\delta_{\varepsilon\varepsilon^{\prime}}\delta_{\nu\nu^{\prime}},~{}(z_{a}|z_{\varepsilon,\nu})=0,~{}(z_{a}|z_{b})=i\Omega_{ab}.$ (122)
The quantum dynamics of the field operator is again described by the time-independent Heisenberg equation of motion (62). However, when working in the Gibbs picture (63), we now expand the field operator around the wavefunction of the spontaneous Floquet state, $\hat{\Psi}^{\prime}(\mathbf{x},t)=\Psi_{0}(\mathbf{x},t)+\hat{\varphi}(\mathbf{x},t)$, arriving at the time-dependent BdGG equations $i\partial_{t}\hat{\Phi}=M_{0}(t)\hat{\Phi}$. Using the complete Floquet spectrum, we can expand the quantum fluctuations of the field operator as
@@
$+$ $\sum_{\varepsilon,\nu}\hat{\gamma}_{\varepsilon,\nu}(t)z_{\varepsilon,\nu}(\mathbf{x},t)+\hat{\gamma}^{\dagger}_{\varepsilon,\nu}(t)\bar{z}_{\varepsilon,\nu}(\mathbf{x},t).$
In general, if $z_{n}(\mathbf{x},t)$ is a generic time-dependent spinor, its related amplitude operator $\hat{\gamma}_{n}(t)\equiv(z_{n}(t)|\hat{\Phi}(t))$ still satisfies $[\hat{\gamma}_{n},\hat{\gamma}^{\dagger}_{m}]=(z_{n}(t)|z_{m}(t))$. After taking into account the time dependence of $z_{n}(\mathbf{x},t)$, the equation of motion for $\hat{\gamma}_{n}(t)$ is simply
$i\partial_{t}\hat{\gamma}_{n}(t)=([M_{0}(t)-i\partial_{t}]z_{n}|\hat{\Phi})$ (124)
This results in the same time dependence for the quantum amplitudes as in the stationary case, Eq. (III.4), namely:
$\hat{\gamma}_{\varepsilon,\nu}(t)=\hat{\gamma}_{\varepsilon,\nu}e^{-i\varepsilon t},\hat{P}^{A}(t)=\hat{P}^{A},\hat{X}^{\alpha}(t)=\hat{X}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}^{B}t.$ (125)
The only problem with this formalism is that the role of the Lagrange multiplier associated with time is not clear since $\lambda^{t}=1$ is fixed by the Schrödinger equation. This is also seen from the fact that the associated charge is the generalized Gibbs Hamiltonian $K$, which should naively satisfy $\partial_{K}E=1$. The corresponding momentum $\hat{P}^{K}=-i(z_{t}|\hat{\Phi})$ describes the quantum fluctuations of $K$,
$\hat{K}=\hat{H}-\lambda^{\alpha}\hat{Q}_{\alpha}\simeq K+\hat{P}^{K}.$ (126)
However, since the preservation of the commutation relations implies $\partial_{A}\lambda^{t}=0=\partial_{K}\lambda^{\alpha}$, Eq. (82), the momentum $\hat{P}^{K}$ does not lead to any dynamics, which is contradictory as fluctuations in $K$ should indeed affect it. Moreover, we lack a satisfying thermodynamical description since now the energy is also an independent variable. All of this suggests that we need to extend our formalism to attain the complete picture.
@@
VI.1 $(t,\phi)$ formalism
In order to correctly derive the dynamics of the temporal FNG mode, we develop the $(t,\phi)$ formalism, denoted in this way due to its analogy with the more usual $(t,t^{\prime})$ formalism Peskin and Moiseyev (1993); Heimsoth et al. (2012). Here, $\phi\in[0,2\pi)$ is the angular variable associated to the periodic motion of the spontaneous Floquet state, $\phi=\omega t^{\prime}$, and $t$ is a proper time that accounts for the nonperiodic dynamics. As we will see, the use of $\phi$ instead of $t^{\prime}$ is critical.
@@ VI.1 $(t,\phi)$ formalism
$L=\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\left(\int\mathrm{d}\mathbf{x}~{}[i\Psi^{*}\partial_{t}\Psi+i\Psi^{*}\omega\partial_{\phi}\Psi]-H\right),$ (127)
where $H$ does not explicitly depend on $\phi$ as the Hamiltonian is time-independent. The case of a conventional, driven Floquet system is simply accounted by a $\phi$-dependent periodic Hamiltonian, $H(\phi+2\pi)=H(\phi)$.
The resulting equation of motion for $\Psi(\mathbf{x},\phi,t)$ is the classical version of Eq. (III) with the substitution
@@
$\mathbf{Q}=\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi~{}\mathbf{Q}(\phi)=\braket{\Psi}{\mathbf{T}}{\Psi},$ (129)
@@
$\braket{\chi}{\Psi}$$\equiv$ $\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\int\mathrm{d}\mathbf{x}~{}\chi^{*}(\mathbf{x},\phi)\Psi(\mathbf{x},\phi)$
$=$ $\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi~{}\braket{\chi(\phi)}{\Psi(\phi)}$
@@
In addition, we now have an extra conserved charge $Q_{\phi}=F$, which we will denote as the Floquet charge, resulting from the invariance under phase translations $\phi\to\phi+\phi_{0}$,
$F=\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\int\mathrm{d}\mathbf{x}~{}i\Psi^{*}\partial_{\phi}\Psi=\braket{\Psi}{T_{\phi}}{\Psi},$ (131)
with $T_{\phi}=i\partial_{\phi}$ the generator of phase translations. The operator $T_{\phi}$ is Hermitian under the extended scalar product (VI.1), and commutes with the rest of generators, $[T_{\phi},\mathbf{T}]=0$. Finally, invariance under time translations $t\to t+t_{0}$ results in the conservation of the Floquet energy
@@
We note that the $(t,\phi)$ formalism can be applied not only to both spontaneous and conventional Floquet systems, but to any many-body system governed by a second-quantization Hamiltonian of the general form (III), including both bosons and fermions, since the dynamics of the field operator then admits a Lagrangian description, as discussed at the beginning of Sec. III.
VI.2 Spontaneous Floquet states in the generalized Gibbs ensemble: Floquet thermodynamics
We invoke again the generalized Gibbs ensemble to describe spontaneous symmetry breaking within the $(t,\phi)$ formalism, replacing the Hamiltonian $H$ by the generalized Gibbs Hamiltonian $K$ in Eq. (127), in analogy with the constrained Lagrangian (20). As a result, the Floquet energy $I$ is replaced by the generalized Floquet-Gibbs energy
@@ VI.2 Spontaneous Floquet states in the generalized Gibbs ensemble: Floquet thermodynamics
which is the analogue of the generalized Gibbs energy $K$ for Floquet systems. Its expression allows us to identify $\lambda^{\phi}=\omega$ as the Lagrange multiplier of the Floquet charge.
In the case of a condensate, the generalized Floquet-Gibbs energy $\Lambda$ is the thermodynamic potential that is extremized by stationary solutions $\Psi_{0}(\mathbf{x},\phi)$, satisfying the Floquet-GPG (FGPG) equation
$00=\frac{\delta\Lambda}{\delta\Psi^{*}}\Longrightarrow\Lambda_{GP}\Psi_{0}\equiv[K_{GP}-i\omega\partial_{\phi}]\Psi_{0}=$ 0 . (134)
When we insert the FGPG wavefunction in the actual time-dependent GP equation (22) by taking $\Psi(\mathbf{x},t=0)=\Psi_{0}(\mathbf{x},\phi_{0})$, we find
$\Psi(\mathbf{x},t)=e^{-i\lambda^{\alpha}T_{\alpha}t}\Psi_{0}(\mathbf{x},\phi_{0})=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\Psi_{0}(\mathbf{x},\phi_{0}+\omega t).$ (135)
Thus, Floquet states are described by stationary FGPG solutions with the phase fixed to $\phi=\phi_{0}$, which behaves as a global time origin $t_{0}=\phi_{0}/\omega$. In a spontaneous Floquet state, this time origin can be chosen freely due to the time-independence of the Hamiltonian. In a conventional Floquet state, $t_{0}$ is fixed by the external driving, as discussed after Eq. (128). On the other hand, as expected from thermodynamic considerations, $\omega$ plays the predicted role for the Lagrange multiplier associated to the Floquet charge, acting as the velocity of phase translations. In fact, the definition modulo $\omega$ of the chemical potential $\mu$ for Floquet states can be also explained from Thermodynamics as follows. The transformation $\Psi_{0}(x,\phi)\to\Psi_{0}(x,\phi)e^{-in\phi}$, $n\in\mathbb{Z}$, also yields a solution of the FGPG equation. This leads to $F\to F+nN$, so $\Lambda\to\Lambda-n\omega N$, and hence $\lambda^{\theta}=\mu$ changes as $\mu\to\mu+n\omega$.
Another hallmark of spontaneous Floquet states is that they conserve energy, $E=H=\Lambda+\lambda^{\alpha}Q_{\alpha}$, since $\{\Lambda,Q_{\alpha}\}$ are conserved charges; the argument fails for driven Floquet systems because there $F$ is not conserved. Following Eqs. (III.1)-(32), it can be shown that each Lagrange multiplier can be obtained from the energy as $\partial_{A}E=\lambda_{A}$, including $\partial_{F}E=\lambda_{F}=\omega$. Moreover, if we now define the local pressure $p(\mathbf{x})$ as the angular average
$p(\mathbf{x})\equiv\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi~{}\frac{|\Psi_{0}(\mathbf{x},\phi)|^{4}}{2},$ (136)
@@
and the Gibbs-Duhem relation (32) for spontaneous Floquet states.
The upshot of the above discussion is that Floquet states allow for a thermodynamical description completely analogous to that of stationary states, which we will denote as Floquet thermodynamics. Indeed, the first principle of Thermodynamics for spontaneous Floquet states (137) suggests a fundamental distinction between spontaneous and driven Floquet systems: spontaneous systems conserve the charge $F$ and the frequency $\omega$ is determined by the equation of state, while driven systems externally impose the frequency $\omega$ so then $F$ is not conserved. Thus, we can regard spontaneous Floquet systems as “isofloquetic”, where the macroscopic Floquet charge is conserved, in analogy with isentropic or isochoric systems, where the entropy or the volume is conserved. Conversely, conventional Floquet systems are “isoperiodic”, where the external driving fixes the oscillation period, in analogy with isothermal or isobaric systems, where the external environment fixes the temperature or the pressure. Based on this analogy, the Floquet energy $I=E-\omega F$ can be identified as the Floquet enthalpy, the natural thermodynamic potential for isoperiodic systems, where it is conserved and satisfies
@@
Physically, the Floquet charge accounts for the independence of the energy, allowing for the emergence of the temporal FNG mode. This is because the stationarity condition for solutions of the Lagrangian (20) imposes a constraint from where the energy becomes a function of the remaining charges, $E=E(\mathbf{Q})$. As argued at the end of Sec. IV, this dependence implies that time-translation symmetry is not spontaneously broken by itself but rather in terms of other symmetries. However, for spontaneous Floquet states, energy becomes an independent charge, and the thermodynamic magnitude whose natural variables are the energy and the other conserved charges is precisely the Floquet charge, $F=F(E,\mathbf{Q})$. This is translated into the emergence of a genuine, independent temporal FNG mode. Remarkably, energy conservation implies that spontaneous Floquet states do not experience heating, in contrast to conventional Floquet systems.
VI.3 Quantization
The $(t,\phi)$ formalism emerges in the quantum theory when considering once more the time-independent Heisenberg equation (62). We now perform a unitary transformation to the field operator along the lines of Eq. (135),
$\hat{\Psi}(\mathbf{x},t)=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\hat{\Psi}(\mathbf{x},\phi+\omega t)=e^{-i\lambda^{\alpha}T_{\alpha}t}\hat{\Psi}(\mathbf{x},\phi,t).$ (139)
The resulting equation of motion for the $(t,\phi)$ field operator reads

### v4

@@ 6 Simultaneous symmetry breaking in spontaneous Floquet states
of the constrained Lagrangian (19), where $\{K,\mathbf{Q}\}$ are still a set of conserved charges. The resulting equation of motion is
@@
When inserted into the usual time-dependent GP equation (21), taking $\Psi(\mathbf{x},0)=\Psi_{0}(\mathbf{x},0)$ yields
@@
$M_{0}(t)-i\partial_{t}]z_{\alpha}=0,\leavevmode\nobreak\ [M_{0}(t)-i\partial_{t}]z_{A}=i(\partial_{A}\lambda^{\beta})z_{\beta},$ (125)
@@
$M_{0}(t)-i\partial_{t}]z_{n}=\varepsilon_{n}z_{n},$ (126)
with $z_{n}(\mathbf{x},t+T)=z_{n}(\mathbf{x},t)$ and $\varepsilon_{n}$ the quasi-energy (defined modulo $\omega=2\pi/T$). With the help from the conventional Floquet theory and from the results of Secs. 3.2, 3.3, it is shown that the orthogonality relations (3.3) still hold:
$(z_{n}|z_{n^{\prime}})=\delta_{nn^{\prime}},\leavevmode\nobreak\ (z_{a}|z_{n})=0,\leavevmode\nobreak\ (z_{a}|z_{b})=i\Omega_{ab}.$ (127)
The dynamics of the field operator is again described by the time-independent Heisenberg equation of motion (66). However, when working in the Gibbs picture (65), we now expand the field operator around the wavefunction of the spontaneous Floquet state, $\hat{\Psi}^{\prime}(\mathbf{x},t)=\Psi_{0}(\mathbf{x},t)+\hat{\varphi}(\mathbf{x},t)$, arriving at the time-dependent BdGG equations $i\partial_{t}\hat{\Phi}=M_{0}(t)\hat{\Phi}$. Using the complete Floquet spectrum, we can express the quantum fluctuations of the field operator as
@@
$+\sum_{n}\hat{\gamma}_{n}(t)z_{n}(\mathbf{x},t)+\hat{\gamma}^{\dagger}_{n}(t)\bar{z}_{n}(\mathbf{x},t).$ (128)
In general, for any time-dependent spinor $z(\mathbf{x},t)$, we can define an amplitude operator as $\hat{\gamma}(t)\equiv(z(t)|\hat{\Phi}(t))$, satisfying $\hat{\gamma},\hat{\gamma}^{\prime\dagger}]=(z(t)|z^{\prime}(t))$, with $\hat{\gamma}^{\prime}(t)=(z^{\prime}(t)|\hat{\Phi}(t))$. After taking into account the time dependence of $z(\mathbf{x},t)$, the equation of motion for $\hat{\gamma}(t)$ is simply
[…]
Applying this relation to each mode results in the same formal time dependence for the quantum amplitudes as in the stationary case, Eq. (3.4), namely:
$\hat{\gamma}_{n}(t)$ $=$ $\hat{\gamma}_{n}e^{-i\varepsilon_{n}t},$
$\hat{P}^{A}(t)$ $=$ $\hat{P}^{A},$ (130)
$\hat{X}^{\alpha}(t)$ $=$ $\hat{X}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}^{B}t.$
[…]
$\hat{K}=\hat{H}-\boldsymbol{\lambda}\cdot\mathbf{\hat{Q}}\simeq K+\hat{P}^{K}.$ (131)
However, since the preservation of the commutation relations implies $\partial_{A}\lambda^{t}=0=\partial_{K}\lambda^{\alpha}$, Eq. (86), the momentum $\hat{P}^{K}$ does not lead to any dynamics, which is contradictory as fluctuations in $K$ should indeed affect the system. Moreover, we lack a satisfying thermodynamical description because now the energy is also an independent variable. All of this suggests that we need to extend our formalism to attain the complete picture.
@@
6.1 $(t,\phi)$ formalism
In order to correctly derive the dynamics of the temporal FNG mode, we develop the $(t,\phi)$ formalism, denoted in this way due to its analogy with the more usual $(t,t^{\prime})$ formalism [58, 59]. Here, $\phi\in[0,2\pi)$ is the angular variable associated to the periodic motion of the spontaneous Floquet state, $\phi=\omega t^{\prime}$, and $t$ is a proper time that accounts for the nonperiodic dynamics. As we will see, the use of $\phi$ instead of $t^{\prime}$ is critical.
@@ 6.1 $(t,\phi)$ formalism
$L=\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\left(\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ [i\Psi^{\dagger}\partial_{t}\Psi+i\Psi^{\dagger}\omega\partial_{\phi}\Psi]-H\right),$ (132)
where $H$ does not explicitly depend on $\phi$ as the Hamiltonian is time-independent. The case of a conventional, driven Floquet system is simply accounted for by a $\phi$-dependent periodic Hamiltonian, $H(\phi+2\pi)=H(\phi)$.
The resulting equation of motion for $\Psi(\mathbf{x},\phi,t)$ is the classical version of Eq. (3) with the substitution
@@
$\mathbf{Q}=\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\leavevmode\nobreak\ \mathbf{Q}(\phi)=\braket{\Psi\|\mathbf{T}\|\Psi},$ (134)
@@
$\braket{\chi\|\Psi}$ $\equiv$ $\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ \chi^{\dagger}(\mathbf{x},\phi)\Psi(\mathbf{x},\phi)$ (135)
$=$ $\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\leavevmode\nobreak\ \braket{\chi(\phi)|\Psi(\phi)}$
@@
In addition, we now have an extra conserved charge $Q_{\phi}=F$, which we will denote as the Floquet charge, resulting from invariance under phase translations $\phi\to\phi+\phi_{0}$:
$F=\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\int\mathrm{d}\mathbf{x}\leavevmode\nobreak\ i\Psi^{\dagger}\partial_{\phi}\Psi=\braket{\Psi\|T_{\phi}\|\Psi},$ (136)
with $T_{\phi}=i\partial_{\phi}$ the generator of phase translations. The operator $T_{\phi}$ is Hermitian under the extended scalar product (135), and commutes with the rest of generators, $T_{\phi},\mathbf{T}]=0$.
Finally, invariance under time translations $t\to t+t_{0}$ results in the conservation of the Floquet enthalpy [see Eq. (143) and ensuing discussion]
@@
We note that the $(t,\phi)$ formalism can be applied to spontaneous and conventional Floquet systems in any many-body system governed by a second-quantization Hamiltonian of the general form (8), including both bosons and fermions, since the dynamics of the field operator then admits a Lagrangian description, as discussed at the beginning of Sec. 3.
6.2 Spontaneous Floquet states in the generalized Gibbs ensemble: Floquet thermodynamics
We invoke again the generalized Gibbs ensemble to describe spontaneous symmetry breaking within the $(t,\phi)$ formalism, replacing the Hamiltonian $H$ by the generalized Gibbs Hamiltonian $K$ in Eq. (132), in analogy with the constrained Lagrangian (19). As a result, the Floquet enthalpy $I$ is replaced by the generalized Floquet-Gibbs energy
@@ 6.2 Spontaneous Floquet states in the generalized Gibbs ensemble: Floquet thermodynamics
which is the analogue of the generalized Gibbs energy for Floquet systems. The expression $\lambda^{\alpha}Q_{\alpha}\equiv\boldsymbol{\lambda}\cdot\mathbf{Q}+\omega F$ allows us to identify $\lambda^{\phi}=\omega$ as the Lagrange multiplier of the Floquet charge.
In the case of a condensate, the generalized Floquet-Gibbs energy $\Lambda$ is the thermodynamic potential that is extremized by stationary solutions $\Psi_{0}(\mathbf{x},\phi)$, satisfying the Floquet-GPG equation
$0=\frac{\delta\Lambda}{\delta\Psi^{*}}\Longrightarrow\Lambda_{GP}\Psi_{0}\equiv[K_{GP}-i\omega\partial_{\phi}]\Psi_{0}=0.$ (139)
When we insert this wavefunction into the actual time-dependent GP equation (21) by taking $\Psi(\mathbf{x},t=0)=\Psi_{0}(\mathbf{x},\phi_{0})$, we find
$\Psi(\mathbf{x},t)$ $=$ $e^{-i\lambda^{\alpha}T_{\alpha}t}\Psi_{0}(\mathbf{x},\phi_{0})$
$=$ $e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\Psi_{0}(\mathbf{x},\phi_{0}+\omega t).$
Thus, Floquet states are described by stationary solutions with $\phi=\phi_{0}$, which behaves as a global time origin $t_{0}=\phi_{0}/\omega$ in the actual GP solution. In a spontaneous Floquet state, this time origin can be chosen freely due to the time-independence of the Hamiltonian. In a conventional Floquet state, $t_{0}$ is fixed by the external driving, as discussed after Eq. (133). On the other hand, as expected from thermodynamic considerations, $\omega$ plays the predicted role for the Lagrange multiplier associated to the Floquet charge, acting as the velocity of phase translations. In fact, the definition modulo $\omega$ of the chemical potential $\mu$ for Floquet states can be also explained from Thermodynamics as follows. The transformation $\Psi_{0}(\mathbf{x},\phi)\to\Psi_{0}(\mathbf{x},\phi)e^{-in\phi}$, $n\in\mathbb{Z}$, also yields a solution of the Floquet-GPG equation. This leads to $F\to F+nN$, so $\Lambda\to\Lambda-n\omega N$, and hence $\lambda^{\theta}=\mu$ changes as $\mu\to\mu+n\omega$.
Another hallmark of spontaneous Floquet states is that they conserve energy, $E=\Lambda+\lambda^{\alpha}Q_{\alpha}$, since $\{\Lambda,Q_{\alpha}\}$ are conserved charges; the argument fails for driven Floquet systems because there $F$ is not conserved. Following Eqs. (3.1)-(31), it can be shown that each Lagrange multiplier can be obtained from the energy as $\partial_{A}E=\lambda_{A}$, including $\partial_{F}E=\lambda_{F}=\omega$. Moreover, if we now define the local pressure $p(\mathbf{x})$ as the angular average
$p(\mathbf{x})\equiv\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\leavevmode\nobreak\ \frac{|\Psi_{0}(\mathbf{x},\phi)|^{4}}{2},$ (141)
@@
and of the Gibbs-Duhem relation (31) for spontaneous Floquet states.
The upshot of the above discussion is that Floquet states allow for a thermodynamical description completely analogous to that of stationary states, which we denote as Floquet thermodynamics. Indeed, the first principle of Thermodynamics for spontaneous Floquet states (142) suggests a fundamental distinction between spontaneous and driven Floquet systems: spontaneous systems conserve the charge $F$ and the frequency $\omega$ is determined by the equation of state, while driven systems externally impose the frequency $\omega$ so then $F$ is not conserved. Thus, we can regard spontaneous Floquet systems as “isofloquetic”, where the macroscopic Floquet charge is conserved, in analogy with isentropic or isochoric systems, where the entropy or the volume is conserved. Conversely, conventional Floquet systems are “isoperiodic”, where the external driving fixes the oscillation period, in analogy with isothermal or isobaric systems, where the external environment fixes the temperature or the pressure. Based on this analogy, $I=E-\omega F$ can be identified as the Floquet enthalpy, the natural thermodynamic potential for isoperiodic systems, where it is conserved and satisfies
@@
Physically, the Floquet charge accounts for the independence of the energy, allowing for the emergence of the temporal FNG mode. This is because the stationarity condition for solutions of the Lagrangian (19) imposes a constraint from where the energy becomes a function of the remaining charges, $E=E(\mathbf{Q})$. As argued at the end of Sec. 4, this dependence implies that time-translation symmetry is not spontaneously broken by itself but rather in terms of other symmetries. However, for spontaneous Floquet states, energy becomes an independent charge, and the thermodynamic magnitude whose natural variables are the energy and the other conserved charges is precisely the Floquet charge, $F=F(E,\mathbf{Q})$. This is translated into the emergence of a genuine, independent temporal FNG mode.
6.3 Quantization
The $(t,\phi)$ formalism arises in the quantum theory when considering once more the time-independent Heisenberg equation (66). In order to work in the rotating frame, we perform a unitary transformation to the field operator along the lines of Eq. (6.2),
$\hat{\Psi}(\mathbf{x},t)=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\hat{\Psi}(\mathbf{x},\phi+\omega t,t)=e^{-i\lambda^{\alpha}T_{\alpha}t}\hat{\Psi}(\mathbf{x},\phi,t).$ (144)
In the specific case of a condensate, the resulting equation of motion for the field operator in the $(t,\phi)$ formalism reads